# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

### v3

@@ 5.2 Q1: Specialised Policy Learning
Learning Diverse Behaviour (Dispersion) Figures 5(a) and 5(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig. 19), while their NoPS counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy, corroborating standard FuPS limitations to learn diverse behaviour. In contrast, HyperMARL (both linear and MLP variants) ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively enable agent specialisation. SND policy diversity measurements (Fig. 5(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
@@
Adaptability (Navigation). Navigation tasks [5] assess adaptability to homogeneous, heterogeneous, and mixed goals (some agents have the same goals, others different). We compare HyperMARL with baselines including DiCo [6]. While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 12 and 13).
@@
Figure 7: Average Reward (IQM, 95% CI) in Navigation for IPPO Variants. HyperMARL adapts robustly across goal configurations—(a) shared, (b) unique, and (c) mixed. Both linear and MLP versions consistently match or outperform IPPO baselines and DiCo, with the margin widening as the number of agents grows. Sample-efficiency curves appear in App. H.4.
@@
Figure 8: IQM and 95% CI of mean win rate in SMAX. Performance of FuPS Recurrent IPPO and MAPPO and HyperMARL (MLP) on SMAX. HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency. Inverval estimates in App. 22.
@@ 5.3 Q2: Effectiveness in Homogeneous Tasks
SMAX. Finally, we evaluate HyperMARL (MLP) on SMAX, where recurrent FuPS is the established baseline [36, 48, 16]. Figure 8 shows while some FuPS variants might exhibit marginally faster initial convergence on simpler maps, HyperMARL achieves comparable final performance on all maps, using the same GRU backbone for partial observability. These results highlights two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge homogenous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig. 7(a))), even with large observation spaces and many agents.
@@
6 Ablations
We ablate two critical components of HyperMARL, gradient decoupling(Sec. 4.3) and initialisation scaling (Sec. 4.2) on the complex Humanoid-v2 and simpler Dispersion tasks. Removing gradient decoupling (HyperMARL w/o GD), forcing joint observation/embedding processing, universally degrades performance (Figure 9). Omitting initialisation scaling (HyperMARL w/o RF (reset fan in/out), which aligns initial generated weights with conventional networks, reveals task-dependent effects: it is crucial for the 17-agent Humanoid, echoing prior findings on hypernetwork initialisation [8], but has minor impact in Dispersion. Thus, while principled initialisation becomes vital with increasing complexity, gradient decoupling is fundamentally essential across tested scenarios.
@@ 6 Ablations
Figure 9: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialization (w/o RF) across environments. Gradient decoupling (a,b) is consistently critical across both environments, while initialisation scaling (c,d) shows greater importance in the complex Humanoid task but less impact in the simpler Dispersion environment.
@@ 7 Related Work
Hypernetworks in RL and MARL. Hypernetworks are effective in single-agent RL for meta-learning, multi-task learning, and continual learning [3, 4, 37, 20]. In MARL, QMIX [35] used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard GRU. Parallel work, CASH [15], conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces. In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
Variants of Parameter Sharing. Selective Parameter Sharing (SePS) [10] shares weights between similar groups of agents, identified via trajectory clustering. Pruning methods [23, 26] split a single network into agent-specific subnetworks.
Learning Diverse Policies. Shared parameters often limit policy diversity [10, 23, 16, 25]. Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories [25], (2) role-based methods [43, 44], or (3) methods that use structural modifications or constraints to induce diversity in agent policies [23, 6, 26]. Outside FuPS/NoPS, HAPPO [49] uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
@@ 8 Conclusion
We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent parameters without modifying the standard RL objective or requiring preset diversity levels. Our results show HyperMARL can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents. We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients. Overall, these findings establish HyperMARL as a promising architecture for diverse MARL tasks. We discuss limitations in App. A, most notably parameter count, which can be remedied by parameter-efficient hypernetworks (e.g., chunked variants [42, 9]).
@@ D.2 Finding a Suitable Distance Function for Policy Diversity
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 10. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) [13], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
@@
[…]
@@
Table 3: Measure Policy Diversity
@@ E.1 Specialisation and Synchronisation Games Description
Specialisation Game. Agents are encouraged to choose distinct actions. In the simplest setting, it is a two-player matrix game where each agent selects between two actions ($A$ or $B$). As shown in Figure 2(a), agents receive a payoff of $1.0$ when their actions differ (creating two pure Nash equilibria on the anti-diagonal) and $0.5$ when they match. This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies. There is also a symmetric mixed-strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
@@
$N$-Agent Extension. Both games naturally scale to $n$ agents and $n$ possible actions. In Specialisation, unique actions receive a payoff of $1.0$, while selecting the same action receives payoffs of $\frac{1}{k}$, where $k$ is the number of agents choosing that action. In contrast, in Synchronisation, agents receive maximum payoffs ($1.0$) only when all actions match. For partial coordination, rewards follow a hyperbolic scale, $\frac{1}{n-k+1}$, encouraging agents to align their choices. Visualisations in Figure 2 and detailed reward profiles appear in Figure 12.
@@ E.2 General-$n$ Payoff Definitions
Temporal Extension. To model sequential decision-making, we extend each normal-form game into a repeated Markov game, where the state at time $t$ is the joint action at time $t-1$. At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff. This repeated setup isolates how agents adapt based on past joint actions, exposing temporal patterns of specialisation and coordination.
@@ E.2.2 Synchronisation Game
Figure 12: Reward profiles for the Specialisation (blue) and Synchronisation (orange) games with $n=5$ agents. In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared. In the Synchronisation game, payoffs follow a hyperbolic scale $1/(n-k+1)$, reaching maximum only under full consensus, thereby incentivising alignment.
@@ F.2 Variance of the HyperMARL Gradient Estimator
Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory-induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
@@ G.1 Environments
SePS [10] Selective Parameter Sharing Requires pretraining phase, which extends beyond the scope of our current study focused on end-to-end learning approaches.

### v4

@@ 5.2 Q1: Specialised Policy Learning
Learning Diverse Behaviour (Dispersion) Figures 4(a) and 4(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig. 19), while their NoPS counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy, corroborating standard FuPS limitations to learn diverse behaviour. In contrast, HyperMARL (both linear and MLP variants) ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively enable agent specialisation. SND policy diversity measurements (Fig. 4(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
@@
Adaptability (Navigation). Navigation tasks [5] assess adaptability to homogeneous, heterogeneous, and mixed goals (some agents have the same goals, others different). We compare HyperMARL with baselines including DiCo [6]. While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 15 and 16).
@@
Figure 6: Average Reward (IQM, 95% CI) in Navigation for IPPO Variants. HyperMARL adapts robustly across goal configurations—(a) shared, (b) unique, and (c) mixed. Both linear and MLP versions consistently match or outperform IPPO baselines and DiCo, with the margin widening as the number of agents grows. Sample-efficiency curves appear in App. H.5.
@@
Figure 7: IQM and 95% CI of mean win rate in SMAX. Performance of FuPS Recurrent IPPO and MAPPO and HyperMARL (MLP) on SMAX. HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency. Interval estimates in Fig. 22 in App. H.6.
@@ 5.3 Q2: Effectiveness in Homogeneous Tasks
SMAX. Finally, we evaluate HyperMARL (MLP) on SMAX, where recurrent FuPS is the established baseline [41, 53, 17]. Figure 7 shows while some FuPS variants might exhibit marginally faster initial convergence on simpler maps, HyperMARL achieves comparable final performance on all maps, using the same GRU backbone for partial observability. These results highlight two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge to homogeneous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig. 6(a))), even with large observation spaces and many agents.
@@
6 Ablations and Embedding Analysis
6.1 Ablations: Gradient Decoupling and Initialisation Scaling
We ablate two critical components of HyperMARL: gradient decoupling (Sec. 4.3) and initialisation scaling (Sec. 4.2). In HyperMARL w/o GD, the hypernetwork is conditioned on $o_{t},e^{i}_{t}]$, coupling observation and agent-ID gradients. In HyperMARL w/o RF, we remove the reset fan-in/out scaling that aligns the scale of generated parameters $(\theta^{i},\phi^{i})$ with standard initialisers.
Gradient decoupling is essential; initialisation scaling grows with complexity. Figure 8 shows that removing GD consistently degrades performance across both Humanoid-v2 (17 agents) and Dispersion, showing that GD is an essential component of HyperMARL. Removing RF reveals a task-dependent effect: it is critical on Humanoid-v2, consistent with hypernetwork initialisation results [8], but has a minor impact on Dispersion, indicating that principled initialisation becomes more vital with increased complexity. We provide additional ablations in App. H.7.
@@ 6.1 Ablations: Gradient Decoupling and Initialisation Scaling
Figure 8: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialisation (w/o RF) across environments. Gradient decoupling (a,b) is consistently critical across both environments, while initialisation scaling (c,d) shows greater importance in the complex Humanoid task but less impact in the simpler Dispersion environment.
Figure 9: Embedding similarity reflects task demands. Mean pairwise cosine distance (dashed line = 1.0). Same goal: contraction ($0.882\pm 0.042$). Different goals: near-orthogonal ($1.010\pm 0.017$). Embeddings contract when a shared policy is optimal and maintain separation for specialisation.
6.2 Analysis of Learned Agent Embeddings
Recall from 4.2 that each agent $i$ possesses an embedding $e^{i}$, which serves as input to the hypernetworks. For the MLP hypernetworks, these embeddings are learned end-to-end and are orthogonally initialised. Consequently, at initialisation (step 0), the pairwise cosine distance between any two distinct agent embeddings is $1.0$, mirroring the separability of one-hot IDs.
We probe how these embeddings $e^{i}$ adapt in the 4-agent Navigation task (same task as in Figure H.5) under two objectives with identical dynamics: (i) same goal (all agents navigate to a single shared target) and (ii) different goals (each agent navigates to its own unique target). At the end of training, we compute the mean pairwise cosine distance between agents’ embeddings (lower values imply greater similarity) and compare it to the orthogonal baseline of $1.0$. (Cosine distance =1cosine similarity=1-$\text${cosine similarity}; 0 = identical direction, $1$ = orthogonal, $2$ = opposite). Figure 9 shows the per-seed distributions, with the $1.0$ initialisation value as a dashed line.
Agent embeddings adapt to task demands. When behaviour should be homogeneous (same goal), embedding directions become significantly more aligned, the mean pairwise cosine distance contracts to $0.882\pm 0.042$ (one-sample $t$-test vs $1.0$: $p=0.0079$). Conversely, when behaviour must differ (different goals), the embeddings maintain their separation, remaining near their orthogonal initialisation at $1.010\pm 0.017$. These results demonstrate that the hypernetwork actively modulates the agent embeddings based on the task, promoting alignment for homogeneity while preserving separability for specialisation. We also conduct a sensitivity analysis on HyperMARL’s hyperparameters in App F.5 and see that agent embedding size can be an important hyperparameter, one that could correspond to the task’s diversity requirements.
@@ 7 Related Work
Hypernetworks in RL and MARL. Hypernetworks are effective in single-agent RL for meta-learning, multi-task learning, and continual learning [3, 4, 42, 21]. In MARL, QMIX [40] used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard GRU. Parallel work, CASH [16], conditions a hypernetwork on local observations and predefined capability descriptors to target zero-shot generalisation across heterogeneous action spaces. By contrast, we use agent-conditioned hypernetworks in homogeneous action spaces, conditioning only on agent IDs/learned embeddings, and we explicitly decouple agent-conditioned from observation-conditioned gradients – a mechanism absent in CASH – which we find critical for specialisation.
Variants of Parameter Sharing. While Full Parameter Sharing (FuPS) is the most common approach, several other variants exist. Selective Parameter Sharing (SePS) [11] shares weights between similar groups of agents, identified by clustering agent trajectories during a pre-training phase. Pruning methods (SNP-PS, Kaleidoscope) [24, 28] split a single network into agent-specific subnetworks using learned agent masks. AdaPS [27] combines clustering and pruning masks to create shared policies for different groups of agents. Concurrent to our work, GradPS [38] identifies neurons with conflicting gradient updates, clones these neurons, and assigns each clone to a group of agents with low gradient conflict. Unlike these works, HyperMARL does not rely on pre-training (SePS), clustering algorithms requiring a preset number of clusters (SePS, GradPS, AdaPS), intricate pruning hyperparameters (SNP-PS, Kaleidoscope), auxiliary diversity losses (Kaleidoscope), or gradient conflict thresholds (GradPS).
Learning Diverse Policies. Shared parameters often limit policy diversity [11, 24, 17, 26]. Proposed solutions include: maximising mutual information between agent IDs and trajectories [26], role-based methods that assign distinct roles to agents [48, 49], best-response [39] methods and approaches that use structural modifications or constraints to induce diversity [24, 6, 28, 27]. Outside FuPS/NoPS, HAPPO [54] uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours. In contrast to these works, HyperMARL does not alter the learning objective, use sequential updates or require preset diversity levels.
@@ 8 Conclusion
We investigated why standard parameter sharing fails at behavioural diversity, identifying that cross-agent gradient interference is critically exacerbated by coupling agent IDs with observations. We hypothesised that explicitly decoupling these gradients would enable adaptivity without prior complexities, and confirmed this using our HyperMARL approach. Our results show this decoupling enables adaptive behaviours (up to 30 agents) and is linked to reduced policy gradient variance. These findings establish gradient decoupling via HyperMARL as a versatile, principled approach for adaptive MARL. We discuss limitations in App. A, most notably parameter count, which can be remedied by parameter-efficient hypernetworks (e.g., chunked variants [47, 9]).
9 Acknowledgements
We would like to thank Samuel Garcin, Max Tamborski, Dave Abel, Timothy Hospedales, Trevor Mcinroe, Elliot Fosong, and Aris Filos-Ratsikas for fruitful discussions on early versions of this work. We also wish to acknowledge the anonymous reviewers for their constructive comments and feedback that helped strengthen this work.
@@ D.2 Finding a Suitable Distance Function for Policy Diversity
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 10. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) [14], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetric, and it satisfies the triangle inequality.
@@
Wasserstein Distance [46] Continuous* Metric Yes $W(p,q)=\left(\inf_{\gamma\in\Gamma(p,q)}\int_{\mathbb{R}\times\mathbb{R}}|x-y|\,d\gamma(x,y)\right)^{1/p}$
@@
Table 4: Measures for Policy Diversity
@@ E.1 Specialisation and Synchronisation Games Description
Specialisation Game. Agents are encouraged to choose distinct actions. In the simplest setting, it is a two-player matrix game where each agent selects between two actions ($A$ or $B$). As shown in Figure 2(a), agents receive a payoff of $1.0$ when their actions differ (creating two pure Nash equilibria on the anti-diagonal) and $0.5$ when they match. This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies. There is also a symmetric mixed‐strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
@@
NN-Agent Extension. Both games naturally scale to $n$ agents and $n$ possible actions. In Specialisation, unique actions receive a payoff of $1.0$, while selecting the same action receives payoffs of $\frac{1}{k}$, where $k$ is the number of agents choosing that action. In contrast, in Synchronisation, agents receive maximum payoffs ($1.0$) only when all actions match. For partial coordination, rewards follow a hyperbolic scale, $\frac{1}{n-k+1}$, encouraging agents to align their choices. Visualisations in Figure 2 and detailed reward profiles appear in Figure 12.
@@ E.2 General-$n$ Payoff Definitions
Temporal Extension. To model sequential decision-making, we extend each normal‑form game into a repeated Markov game, where the state at time $t$ is the joint action at time t1t-1. At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff. This repeated setup isolates how agents adapt based on past joint actions, exposing temporal patterns of specialisation and coordination.
@@ E.2.2 Synchronisation Game
Figure 12: Reward profiles for the Specialisation (blue) and Synchronisation (orange) games with $n=5$ agents. In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared. In the Synchronisation game, payoffs follow a hyperbolic scale 1/(nk+1)1/(n-k+1), reaching maximum only under full consensus, thereby incentivising alignment.
@@ F.2 Variance of the HyperMARL Gradient Estimator
Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory‐induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
@@ F.4 Speed and Memory Usage
F.5 Sensitivity Analysis
We run a sensitivity analysis on the HyperMARL’s hyperparameters in the 20-agent SMAX scenario with recurrent MAPPO, a challenging setting for optimisation. The results are shown Table 5.
Findings.
  • Similarly to FuPS, and in general, deep RL methods like PPO, the learning rate is an important parameter for performance.
  • Agent embedding size is an important hyperparameter. In these results, a smaller embedding size outperformed a larger one. This could be due to the homogeneous nature of some SMAX tasks, where with smaller embeddings, it could be easier to learn similar agent embeddings and hence similar behaviours. This suggests the embedding size could correspond to the task’s diversity requirements.
  • Other parameters, such as the width, have limited impact beyond a modest size.
Table 5: IQM and 95% CI of mean win rate across 5 seeds for the 20-agent SMAX scenario with recurrent MAPPO. Bold indicates the highest IQM score; * indicates scores whose 95% confidence intervals overlap with the highest score.
Method Embedding Dim Hidden Dims LR 0.0001 LR 0.0003 LR 0.0005
FuPS N/A N/A 0.1067 (0.0513, 0.2128) 0.3053 (0.2041, 0.4769) *0.4213 (0.3623, 0.4800)
HyperMARL 4 16 0.0799 (0.0250, 0.1190) 0.2946 (0.1143, 0.3778) 0.4455 (0.3443, 0.5692)
64 0.1233 (0.0750, 0.2162) 0.3338 (0.2692, 0.4286) *0.3765 (0.3281, 0.4107)
128 0.1147 (0.0270, 0.1842) *0.3787 (0.3509, 0.4603) 0.3327 (0.2931, 0.3548)
64, 64 0.1781 (0.1628, 0.2075) *0.3498 (0.2807, 0.4500) 0.1944 (0.1268, 0.3077)
HyperMARL 64 16 0.1109 (0.0256, 0.2093) 0.3198 (0.2353, 0.3934) *0.3483 (0.2581, 0.4308)
64 0.1360 (0.0303, 0.2821) 0.1191 (0.0517, 0.1964) 0.1155 (0.0794, 0.1500)
128 0.1193 (0.0750, 0.1579) 0.1036 (0.0517, 0.1594) 0.1106 (0.0455, 0.1587)
64, 64 0.1353 (0.0541, 0.1818) 0.1283 (0.0526, 0.2258) 0.0860 (0.0345, 0.1343)
@@ G.1 Environments
Blind-Particle Spread (BPS) [11] Blind-Particle Spread (BPS) is a modified variant of the Multi-Agent Particle Environment (MPE) Simple Spread task [30] with 15 to 30 agents. The environment contains landmarks of different colours, and each agent is assigned a colour. Agents are ’blind’ as they cannot observe their own assigned colour or the colours of other agents. They must infer the correct landmark to navigate towards based only on their observations and rewards. The number of colours thus dictates the number of distinct roles or behaviours the team must learn to successfully complete the task.
@@
SePS [11] Selective Parameter Sharing Although this requires a pre-training phase, it is a strong baseline for parameter sharing approaches. Tested in: Blind-Particle Spread environments, with 15 to 30 agents.